OpenStudy (anonymous):

state the maximum number of turns in the graph of functions f(x)=x^4+2x^2+4??

OpenStudy (anonymous):

\[f(x)=x^4+2x^2+4\] \[f'(x)=4x^3+4x\] \[f''(x)=12x^2+4\] At S.P's (Stationary Points) f'(x)=0 \[4x^3+4x=0\] \[x(x^2+1)=0\] \[x=0\] \[y=4\] \[f''(x)=12(0)^2+4\] \[=4\] \[>0\] Therefore minimum turning point at (0, 4) Therefore there is only 1 turn in the function.